Two methods can be used for producing expansion anchors. Method A costs $85,000 initially and will have a $15,000 salvage value after 3 years. The operation cost with this method will be $32,000 per year. Method B will have a first cost of $125,000, and operating cost of $10000 per year and a $40,000 salvage value after its 3-year life. Using a Future Worth analysis and an interest rate of 8%/yr/yr, which method should you recommend?


please do not use excel
Answer #1

Future worth of option A = -85000 * (F/P, 8%, 3) - 32000 *(F/A, 8%, 3) + 15000

= -85000 * 1.259712 - 32000 * 3.2464 + 15000

= -195960.32

Future worth of option B = -125000 * (F/P, 8%, 3) - 10000 *(F/A, 8%, 3) + 40000

= -125000 * 1.259712 - 10000 * 3.2464 + 40000

= -174928

As Future Cost of option B is less, it should be selected